which equation represents the relationship between x and y
x y
-1 0
0 -1
1 0
3 8
A. y=x+2 B. y=2x+2 C. y=x^2-1 D. y= 2x^2+2
Please help and thank you



Answer by ewatrrr(24785)

Hi,
*Note: the equation must satisfy each and every point given.
.
take one. for ex: (0,-1)
A. y=x+2 no
B. y=2x+2 no
C. y=x^2-1 yes
D. y= 2x^2+2 no
.
C would be the correct answer
.
.
Note had we started with pt(3,8)for ex, we would have needed to eliminate B by trying another point.
B. y=2x+2
C. y=x^2-1